i need help on this algebra problem but i dont understand 8+2(-4x-7)=35.6
Answer by tutorcecilia(2152)

8+2(-4x-7)=35.6 [Use the Order of Operations -- PEMDAS to simplify]
8-8x-14=35.6
-6-8x=35.6 [Solve for the x-term]
-6+6-8x=35.6+6
-8x=41.6
-8x/-8=41.6/-8
x=-5.2
.
Check by plugging x= -5.2 back into the original equation and solve.